Suppose a professor gets 3 visitors per week on average. Let the random variable X be the number of visitors the professor gets during a certain week.

a) What type of distribution does X have?

   ANS: Poisson Distribution

b) What. is the probability mass function, f(x)? What is E(X)? What is V(X)?

c) What is the probability that this teacher receives betweeen 2 and 5 visitors this week?
Answer #1

a? Given ht ne vamabte X Poisson3) nCFor vat(x) E(x21- E(X) = E (1x) (x-1) t ) (E(X)) - 2 and xl 叭 2. 3 = 0.49 2892